# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection

### v1

@@ 8. Conclusions and Future Work
In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection. We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data. Our analysis reveals the efficiency of PEFT methods compared to full fine-tuning, underscoring the potential of PEFT methods to enhance method-level code smell detection and offer valuable insights for optimizing LMs for other software engineering tasks. Additionally, we have constructed high-quality datasets for CC and CM code smells, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
In the next step, we plan to further explore how PEFT performs in cross-language scenarios for code smell detection. Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in code smell refactoring, thereby improving overall code quality and maintainability by providing just-in-time code smell detection and refactoring suggestions.
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.

### v2

@@ 8. Conclusions and Future Work
In this work, we investigated the effectiveness of various PEFT methods on different LMs for code smell detection. We also evaluated how the performance of these PEFT methods varies under different conditions, such as varying hyper-parameter settings and the availability of training data. Our analysis reveals that PEFT methods are more efficient than full fine-tuning and In-Context Learning (ICL) in most cases, underscoring the potential of PEFT methods to enhance code smell detection and offer valuable insights for optimizing LMs for other Software Engineering (SE) tasks. Additionally, we have constructed high-quality datasets for four widely studied code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class, and we provided the replication package of this work online (replicationPackage, ) to facilitate future research and ensure the reproducibility of our study results.
In the next step, we plan to further explore how PEFT performs in cross-language scenarios for code smell detection. Moreover, we plan to investigate the application of PEFT methods for fine-tuning LMs to assist in a wider range of SE tasks, such as bug detection, code completion, and code smell refactoring. These directions will help us better understand the generalizability and practical value of PEFT methods in real-world software development settings.
@@ Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311. The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.